A rocket moves horizontally with a constant velocity ‘u’ at a height ‘L’. A guided missile is fired vertically with a speed ‘v’ when the rocket passes above it. Assuming that the missile always aims at the rocket with the constant speed ‘v’, find the time after which the missile strikes the rocket.

Grade:9

1 Answers

Sher Mohammad IIT Delhi
8 years ago
The plans will strike when the both covers the distance L, the verticl velocity of above plane is, 
v=gt donwards, 
so the relative velocity is gt+v
s= u+at
L=gt+v+gt^2
gt2+gt+v-L=0, solve this qurdatic to get the answer.